All slow machines (except c1auxex) were dead today, so I had to key them all. While I was at it, I also decided to update MC autolocker screen. Kira pointed out that I needed to change the EPCIS input type (in the RTCDS model) to a "binary input", as opposed to an "analog input", which I did. Model recompilation and restart went smooth. I had to go into the epics record manually to change the two choices to "ENABLE" and "DISABLE" as opposed to the default "ON" and "OFF". Anyways, long story short, MC autolocker controls are a bit more intuitive now I think.
A follow-up on the discussion from today's lunch meeting - Rana pointed out that rotation of the fiber in the mount by ~5degrees cannot account for such large power fluctuations. Here is a 3 day trend from my polarization monitoring setup. Assuming the output fiber coupler rotates in its mount by 5 degrees, and assuming the input light is aligned to one of the fiber's special axes, then we expect <1% fluctuation in the power. But the attached trend shows much more drastic variations, more like 25% in the p-polarization (which is what I assume we use for the beat, since the majority of light is in this polarization, both for PSL and EX). I want to say that the periodicity in the power fluctuations is ~12hours, and so this fluctuation is somehow being modulated by the lab temperature, but unfortunately, we don't have the PSL enclosure temperature logged in order to compare coherence.
Steve: your plots look like temperature driven
The "beat length" of the fiber is quoted as <=2.7mm. This means that a linearly polarized beam that is not oriented along one of the special axes of the fiber will be rotated through 180 degrees over 2.7mm of propagation through the fiber. I can't find a number for the coefficient of thermal expansion of the fiber, but if temperature driven fluctuations are changing the length of the fiber by 300um, it would account for ~12% power fluctuation between the two polarizations in the monitoring setup, which is in the ballpark we are seeing...
I've been looking into recovering the seismic BLRMs for the BS Trillium seismometer. It looks like the problem is probably in the anti-aliasing board. There's some heavy stuff sitting on top of it in the rack, so I'll take a look at it later when someone can give me a hand getting it out.
In detail, after verifying that there are signals coming directly out of the seismometer, I tried to inject a signal into the AA board and see it appear in one of the seismometer channels.
It looks like the drift in polarization content in the PSL pickoff is actually much higher than that in the EX pickoff. Note that to prevent the P-pol diode from saturating, I put an ND filter in front of the PD, so the Y axis actually has to be multiplied by 10 to compare power in S and P polarizations. If this drift is because of the input (linear) polarization being poorly matched to one of the fiber's special axes, then we can improve the situation relatively easily. But if the polarization drift is happening as a result of time-varying stress (due to temp. fluctuations, acoustics etc) on the (PM) fiber from the PSL fiber coupler to the BeatMouth, then I think this is a much more challenging problem to solve.
I'll attempt to quantify the contribution (in Hz/rtHz) of beat amplitude RIN to phase tracker output noise, which will tell us how much of a problem this really is and in which frequency bands. In particular, I'm interested in seeing if the excess noise around 100 Hz is because of beat amplitude fluctuations. But on the evidence thus far, I've seen the beat amplitude drift by ~15 dB (over long timescales) on the control room network analyzer, and this drift seems to be dominated by PSL light amplitude fluctuations.
I swapped the EX fiber for the PSL fiber in the polarization monitoring setup. There is a lot more power in this fiber, and one of the PDs was saturated. I should really have put a PBS to cut the power, but I opted for putting an absorptive ND1.0 filter on the PD instead for this test. I want to monitor the stability in this beam and compare it to the EX beam's polarization wandering.
THIS CALCULATION IS WRONG FOR THE OVERCOUPLED CAV.
Mode-matching efficiency of EX green light into the arm cavity is ~70*%, as measured using the visibility.
I wanted to get an estimate for the mode-matching of the EX green beam into the arm cavity. I did the following:
This amount of mode-matching is rather disappointing - using a la mode, the calculated mode-matching efficiency is nearly 100%, but 70% is a far cry from this. The fact that I can't improve this number by either tweaking the steering or by moving the MM lenses around suggests that the estimate of the target arm mode is probably incorrect (the non-gaussianity of the input beam itself is not quantified yet, but I don't believe this input beam can account for 30% mismatch). For the Y-arm, the green REFL DC level is actually higher when locked than when ITMY is misaligned. WTF?? Only explanation I can think of is that the PD is saturated when green is unlocked - but why does the ADC saturate at ~3000cts and not 32000?
This data is almost certainly bogus as the AA box at 1X9 is powered by +/-5VDC and not +/-15VDC. I didn't check but I believe the situation is the same at the Y-end.
3000 cts is ~1V into the ADC. I am going to change the supply voltage to this box (which also reads in ETMX OSEMS) to +/-15V so that we can use the full range of the ADC.
gautam Apr 26 630pm: I re-did the measurement by directly monitoring the REFLDC on a scope, and the situation is not much better. I calculate a MM of 70% into the arm. This is sensitive to the lens positions - while I was working on the EX fiber coupling, I had bumped the lens mounted on a translational stage on the EX table lightly, and I had to move that lens around today in order to recover the GTRX level of 0.5 that I am used to seeing (with arm aligned to maximize IR transmission). So there is definitely room for optimization here.
Attachment #1 shows the drift of the polarization content of the light from EX entering the BeatMouth. Seems rather large (~10%). I'm going to tweak the X end fiber coupling setup a bit to see if this can be improved. This performance is also a good benchmark to compare the PSL IR light polarization drift. I am going to ask Steve to order Thorlabs K6XS, which has a locking screw for the rotational DoF. With this feature, and by installing some HWPs at the input coupling point, we can ensure that we are coupling light into one of the special axes in a much more deterministic way.
changed the setpoint of the EX Seismomter T ctrl servo from 35 to 39 C to see if this helps the stability by decreasing the cooldown time constant.
Today, I decided to check the power coupled into the PSL fiber for the BeatMouth. Surprisingly, it was only 200uW, while I had ~3.15mW going into it in January. Presumably some alignment drifting happened. So I re-aligned the beam into the fiber using the steering mirror immediately before the fiber coupler. I managed to get ~2.9mW in without much effort, and I figured this is sufficient for a first pass, so I didn't try too much more. I then tried making an ALS beat spectrum measurement (arm locked to IMC length using POX, green following the arm using end PDH servo). Surprisingly, the noise performannce was almost as good as the reference! See Attachment #1, in which the red curve is an IR beat (while all others are green beats). The Y arm green beat performance isn't stellar, but one problem at a time. Moreover, the kind of coherence structure between the arm error signal and the ALS beat signal that I reported here was totally absent today.
Upon further investigation, I found that the noise level was actually breathing quite significantly on timescales of minutes. While I was able to successfully keep the TEM00 mode of the PSL beam resonant inside the arm cavity by using the ALS beat frequency as an error signal and MC2 as a frequency actuator, the DC stability was very poor and TRX was wandering around by 50%. So my new hypothesis is that the excess ALS noise is because of one or more of
While I did some work in trying to align the PSL IR pickoff into the fiber along the fast (P-pol) axis, I haven't done anything for the X end pickoff beam. So perhaps the fluctuations in the EX IR power is causing beatnote amplitude fluctuations. In the delay line + phase tracker frequency discriminator, I think RF beatnote amplitude fluctuations can couple into phase noise linearly. For such an apparently important noise source, I can't remeber ever including it in any of the ALS noise budgets.
Before Ph237 today, I decided to use my polarization monitoring setup and check the "RIN" of power in the two polarizations coming out of the fiber on the PSL table. For this purpose, I decided to hijack the Acromag channels used for the PSL diagnostics connector Attachment #2 shows that there is fluctuations at the level of ~10% in the p-polarization. This is the "desired" polarization in that I aligned the PSL beam into the fiber to maximize the power in this polarization. So assuming the power fluctuations in the PSL beam are negligible, this translates to sqrt(10) ~3% fluctuation in the RF beat amplitude. This is at best a conservative estimate, as in reality, there is probably more AM because of the non PM fibers inside the beatmouth.
All of this still doesn't explain the coherence between the measured ALS noise and the arm error signal at 100s of Hz (which presumably can only happen via frequency noise in the PSL).
Another "mystery" - yesterday, while I was working on recovering the Y arm green beat signal on the PSL table, I eventually got a beat signal that was ~20mVpp into 50ohms, which is approximately the same as what I measured when the Y arm ALS performance was "nominal", more than a year ago. But while viewing the Y arm beats (green and IR) simultaneously on an o'scope, I wasn't able to keep both signals synchronised while triggering on one (even though the IR beat frequency was half the green beat frequency). This means there is a huge amount of relative phase noise between the green and IR beats. What (if anything) does this mean? The differential noise between these two beats should be (i) phase noise at the fiber coupler/ inside the fiber itself, and (ii) scatter noise in the green light transmitted through the cavity. Is it "expected" that the relative phase noise between these two signals is so large that we can't view both of them on a common trigger signal on an o'scope? Also - the green mode-matching into the Y arm is abysmal.
Anyways - I'm going to try and tweak the PER and mode-matching into the X end fiber a little and monitor the polarization stability (nothing too invasive for now, eventually, I want to install the new fiber couplers I acquired but for now I'll only change alignment into and rotation of the fiber coupler on the EX table). It would also be interesting to compare my "optimized" PSL drift to the unoptimized EX power drift. So the PSL diagnostic channels will not show any actual PSL diagnostic information until I plug it back in. But I suspect that the EPICS record names and physical channel wiring are wrong anyways - I hooked up my two photodiode signals into what I would believe is the "Diode 1 Power" and "Laser crystal temperature" monitors (as per the schematic), but the signals actually show up for me in "Diode 2 Power" (p-pol) and "Didoe 1 Temperature" (s-pol).
Annoyingly, there is no wiring diagram - on my todo list i guess...
@Steve - could you please take a photo of the EX table and update the wiki? I think the photo we have is a bit dated, the fiber coupler and transmon PDs aren't in it...
Another update. I've changed the on/off button so that it's visible which state it's in. I did that by changing the type of C1:PEM-SEIS-EX_TEMP_SLOWLOOP from ai to bi (I checked the FSS script and copied the entry for the slowloop). Previously, MEDM was giving me an error that it wasn't an ENUM value when I wanted to use a choice button to indicate the value of slowloop, and this solved the issue. I've also added a StripTool button.
Satellite amplifiers labeled with date. Old labels left on.
activities done today - details/plots/data/evidence tomorrow.
An update to the screen. I changed the min/max values for some of the parameters, as well as changing the script so that I could specify the integral gain in terms of 1e-5. I've also added this screen to the PEM tab in the sitemap.
I am working on IMC electronics. IMC is misaligned until further notice.
I think we can scrap the 126MOPA channels since they're associated with the Lightwave NPRO and MOPA. We should add the channels that we need for monitoring the Innolight NPRO from the d-sub connector on its controller.
I made a list of all the physical c1psl channels to get a better idea for how many acromags we need to replace it eventually. There 3123 unit is the one whose failure had prevented c1psl from booting, which is why it was unplugged (elog post 12852), and its channels have been inactive since. Are the 126MOPA channels used for the current mephisto? 126 tells me it's for an old lightwave laser, but I was checking a few and found that they have non-zero, changing values, so they may have been rewired.
It also hosts some virtual channels for the ISS with root C1:PSL-ISS_ defined in iss.db and dc.db, the PSL particle counter with root C1:PEM- defined in PCount.db and a whole lot of PSL status channels defined in pslstatus.db. Transfering these virtual channels to a different machine is almost trivial, but the serial readout of the particle counter would have to find a new home.
Long story short - we need:
I created an MEDM screen for the PID control. In addition, I added a new EPICS channel for the setpoint so that it could be adjusted using the MEDM screen.
Edit: forgot to mention the channel name is C1:PEM-SEIS_EX_TEMP_SETPOINT
Edit #2: the path for the MEDM is /opt/rtcds/caltech/c1/medm/c1pem/C1PEM_SEIS_EX_TCTRL.adl
Earth quake M5.3 2018-04-05 19:29:16UTC Santa Cruz Island, CA
Why is MC2 LR so different from the others???
The previous measurements were made from the coil driver output monitors. To investigate why the MC2 LR coil has less noise than the other coils, I also measured the noise at the output to the coils.
The circuit diagram for the coil driver board is given in D010001 and a picture of the rack is on the 40m wiki here. The coil driver boards are in the upper left quadrant of the rack. The input to the board is the column of LEMOs which are the "Coil Test In" inputs on the schematic. The output monitors are the row of LEMOs to the right of the input LEMOs and are the "FP Coil Volt Mon" outputs on the schematic. The output to the coils "Coil Out" in the schematic are carried through a DB15 connector.
The attachment shows the voltage noise for the MC2 LR coil as well as the UL coil which is similar to all of the other coils measured in the previous measurement. While the LR coil is less noisy than the UL coil as measured at the output monitor, they have the same noise spectrum as measured at the output to the coils themselves. So there must be something wrong with the buffer circuit for the MC2 LR voltage monitor, but the output to the coils themselves is the same as for the other coils.
While Kevin is working on the MC2 electronics chain - we disconnected the output to the optic (DB15 connector on coil driver board). I decided to look at the 'free' freeswinging MC2 OSEM shadow sensor data. Attachment #1 suggests that the suspension eigenmodes are showing up in the shadow sensors, but the 0.8Hz peak seems rather small, especially compared to the result presented in this elog.
Maybe I'll kick all 3 MC optics tonight and let them ringdown overnight, may not be a bad idea to checkup on the health of the MC suspensions/satellite boxes... [MC suspensions were kicked @1207113574]. PSL shutter will remain closed overnight...
Made some changes:
There is also now a StripTool file in the scripts/PEM directory which has appropriate channel names and scales for PID loop tuning. Use this file!
I'm leaving it running over the weekend with K_I = -0.003. There is a StripTool on rossa which you can watch. The code itself is running on a tmux session on megatron. Let's ONLY run this code there until we're satisfied that things are good.
Update Sun Apr 8 00:40:11 2018: Lowered gain by factors of 3 down to -0.0001 Saturday afternoon. Seems like still oscillating a bit, now with a ~4 hour period. Setting it to -3e-5 now. Usually we have a linear feedback loop, but our actuation voltage actually gets squared (P = I^2 R) before being integrated to produce temperature. Wonder if we should think of linearizing the feedback control signal to make the loop act nicer.
Update Sun Apr 8 21:09:48 2018: Set K_I = -1e-5 earlier today. Seems to have stabilized nearly, but temperature swings are still +/- 1 K. Will need to add some proportional feedback (K_P) to increase the loop bandwidth, but system is at least sort of stable now. Probably should start construction of EY,BS systems now.
I have been trying to tune the PID and have managed to descrease the oscillations without saturating the actuator. I'm going to model the system to calculate the exact values of P, I and D in order to get rid of the oscillations altogether. I was going to record the data using Data Viewer, but there seems to be some issue with that, so I'm using StripTool for now.
Kira informed me that she was having trouble accessing past data for her PID tuning tests. Looking at the last day of data, it looks like there are frequent EPICS data dropouts, each up to a few hours. Observations (see Attachment #1 for evidence):
It is difficult to diagnose how long this has been going on for, as once you start pulling longer stretches of data on dataviewer, any "data freezes" are washed out in the extent of data plotted.
Clearly, there is a discrepancy for f>20Hz. Why?
I found all the EPICS channels for the model c1ioo on the FE c1ioo to be blank just now. The realtime model itself seemed to be running fine, judging by the IMC alignment (as the WFS loops seemed to still be running okay). I couldn't find any useful info in demsg but I don't know what I'm looking for. So my guess is that somehow the EPICS process for that model died. Unclear why.
Seems like there was a 5.3 magnitude EQ ~10km from us (though I didn't feel it). All watchdogs were tripped so our mirrors definitely felt it. ITMX is stuck (but all the other optics are damping fine). I tried the usual jiggling of DC bias voltage but ITMX still seems stuck. Probably a good sign that the magnet hasn't come off, but not ideal that I can't shake it free..
edit: after a bit more vigorous shaking, ITMX was freed. I had to move the bias slider by +/-10,000 cts, whereas initially I was trying +/-2000 cts. There is a tendency for the optic to get stuck again once it has been freed (while the optic's free swinging motion damps out), so I had to keep an eye out and as soon as the optic was freed, I re-engaged the damping servos to damp out the optic motion quickly.
I'm probably doing something stupid - but I've not been able to figure this out. In the MC1 and MC3 coil driver filter banks, we have a digital "28HzELP" filter module in FM9. Attachment #1 shows the MC1 filterbanks. In the shown configuration, I would expect the only difference between the "IN1" and "OUT" testpoints to be the transfer function of said ELP filter, after all, it is just a bunch of multiplications by filter coefficients. But yesterday looking at some DTT traces, their shapes looked suspicious. So today, I did the analysis entirely offline (motivation being to rule out DTT weirdness) using scipy's welch. Attachment #2 shows the ASDs of the IN1 and OUT testpoint data (collected for 30s, fft length is set to 2 seconds, and hanning window from scipy is used). I've also plotted the "expected" spectral shape, by loading the sos coefficients from the filter file and using scipy to compute the transfer function.
Code used to generate this plot (and also a datafile to facilitate offline plotting) is attached in the tarball Attachment #3. Note that I am using a function from my Noise Budget repo to read in the Foton filter file...
*ChrisW suggested ruling out spectral leakage. I re-ran the script with (i) 180 seconds of data (ii) fft length of 15 seconds and (iii) blackman-harris window instead of Hanning. Attachment #4 shows similar discrepancy between expectation and measurement...
We measured the MC coil driver noise at the output monitors of the coil driver board with an SR785 in order to further diagnose the excess IMC frequency noise.
Attachments 1 and 2 show the noise for the UL coils of MC3 and MC2 with various combinations of output filters engaged. When the 28 Hz elliptic filter is on, the analog dewhitening filter is off, and vice versa. The effect of the analog low pass filter is visible in MC3, but the effect of the digital low pass filter is swamped by the DAC noise.
We locked the arms and measured the ALS beatnote in each of these filter combinations, but which filters were on did not effect the excess IMC frequency noise. This suggests that the coil drivers are not responsible for the excess noise.
Attachment 2 shows the noise for all five coils on MC1, MC2, and MC3 as well as for ITMY, which is on a different DAC card from the MCs. All filters were on for these measurements.
It's been a while - but today, all slow machines (with the exception of c1auxex) were un-telnetable. c1psl, c1iool0, c1susaux, c1iscaux1, c1iscaux2, c1aux and c1auxey were rebooted. Usual satellite box unplugging was done to avoid ITMX getting stuck.
I did a step response for the loop from 35 degrees to 40 degrees. The PID is not properly tuned, so the signal oscillates. In the graph, the blue curve is the temperature of the can in celcius and the green curve is the heating power in watts. The x-axis is in minutes. Before, the signal was too noisy to do a proper step response, so I placed a 3.3 microF capacitor in parallel with the resistor in my temperature sensor circuit (I'll draw and attach this updated version). This created a 5 Hz low pass filter and the signal is now pretty clean.
I also added in new Epics channels so that we could log the data using Data Viewer. The channels I added were C1:PEM-SEIS_EX_TEMP_MON_CELCIUS and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS. I used 13023 as a guide on how to do this.
Update: the channels work and show data in Data Viewer
Edit: I've attached a photo of the circuit with the capacitor indicated. It is in parallel with the resistor below it. I've attached an updated circuit diagram as well.
The new matching circuit was tested.
f_nominal f_actual response required mod. drivng power
[MHz] [MHz] [mrad/V] [rad] needed [dBm]
9.1 9.1 55 0.22 => 22
118.3 118.2 16 0.01 => 6
45.5 45.4 45 0.28 => 25
24.1 N/A 2.1 0.014 => 27
- 9.1MHz and 118.3MHz: They are just fine.
- 24.1MHz: Definitely better (>x3) than the previous trial to combine 118MHz & 24MHz.
We got about the same modulation with the 50Ohm terminated bare crystal (for the port1).
So, this is sort of the best we can do for the 24.1MHz with the current approach.
The driving power of 27dBm is required at 24.1MHz
- About the 45MHz
- The driving power of 27dBm is required at 24.1MHz
- The maximum driving power with the AM stabilized driver is 23dBm, nominally to say.
- I wonder how we can reduce resistance (and capacitance) of the 45MHz further...?
- I also wonder if the IFO can be locked with reduced modulation (0.28 rad->0.2 rad)
- Can the driver max power be boosted a bit? (i.e. adding an attenuator in the RF power detection path)
We redid the measurement measuring the voltage noise from the REFL PD demod board output monitor with an SR785 with the noise eater on and off. We used a 100x preamp to amplify the signal above the SR785 noise. The SR785 noise floor was measured with the input to the preamp terminated with 50 ohms. The spectra shown have been corrected for the 100x amplification.
This measurement shows no difference with the noise eater on or off.
the noise eater on/off measurements should be done for 0-100 kHz and from the demod board output monitor
I created two new channels today, C1:PEM-SEIS_EX_TEMP_MON_CELCIUS, which turns the output voltage signal into degrees C, and C1:PEM-SEIS_EX_TEMP_CTRL_WATTS, which takes the input voltage from the DAC and turns it into a value of watts. I'm trying to stabilize the temperature at 35 degrees, but it's taking a lot longer than expected. Perhaps we'll need to use different values for P and I and decrease the noise in the sensor, since right now there's about a 10 degree variation between the highest and lowest values.
Can't really figure out what this plot means. We need to see the sensor (in units of deg C) and the control signal (in heating power (W)). The plot should show a few step responses with the PID loop on, so that we can see the loop response time. Please zoom in on the axes so that we can see what's happening.
We closed the loop today and implemented the PID script. I have attached the StripTool graph for an integral value of 0.5 and proportional value of 20. We had some issues getting it to work properly and it would oscillate between some low values of the control voltage. The set point here was -3.20, which corresponds to about a 20 degree increase in temperature. The next step would be to find which values of Kp, Ki, and Kd would work in this case and low pass filter the signal from the temperature sensor, and also create an MEDM screen for easier PID control.
I have been working on the aux beat setup on the PSL table between 9PM-3AM.
This work involved:
- Turning off the main marconi
- Turning off the freq generation unit (incl IMC modulation)
- Closing the PSL shutter
After the work, these were reverted and the IMC and both arms have been locked.
Kevin, Gautam and Arijit
Today we performed the in-loop noise measurements of the MCREFL-PD using the SR785 to ascertain the effect of the Noise Eater on the laser. We took the measurements at the demodulated output channel from the MCREFL-PD. We performed a series of 6 measurements with the Noise Eater ''ON'' and ''OFF''. The first data set is an outlier probably, due to some transient effects. The remaining data sets were recorded in succession with a time interval of 5 minutes each between the Noise Eater in the ''ON'' and ''OFF'' state. We used the calibration factor of 13kHz/Vrms from elog 13696 to convert the V_rms to Hz-scale.
The conclusion is that the NOISE EATER does not have any noticeable effect on the noise measurements.
ALS beat spectrum and also the arm control signal look as they did before. coherence between arm control signals (in POX/POY lock) is high between 10-100Hz, so looks like there is still excess frequency noise in the MC transmitted light. Looking at POX as an OOL sensor with the arm under ALS control shows ~10x the noise at 100 Hz compared to the "nominal" level, consistent with what Koji and I observed ~3weeks ago.
We tried swapping out Marconis. Problem persists. So Marconi is not to blame. I wanted to rule this out as in Jan, Steve and I had installed a 10MHz reference to the rear of the Marconi.
I've been developing an idea for making a direct measurement of the SRC Gouy phase at RF. It's a very different approach from what has been tried before. Prior to attempting this at the sites, I'm interested in making a proof-of-concept measurement demonstrating the technique on the 40m. The finesse of the 40m SRC will be slightly higher than at the sites due to its lower-transmission SRM. Thus if this technique does not work at the 40m, it almost certainly will not work at the sites.
The idea is, with the IFO locked in a signal-recycled Michelson configuration (PRM and both ETMs misaligned), to inject an auxiliary laser from the AS port and measure its reflection from the SRC using one of the pre-OMC pickoff RFPDs. At the sites, this auxiliary beam is provided by the newly-installed squeezer laser. Prior to injection, an AM sideband is imprinted on the auxiliary beam using an AOM and polarizer. The sinusoidal AOM drive signal is provided by a network analyzer, which sweeps in frequency across the MHz band and demodulates the PD signal in-phase to make an RF transfer function measurement. At the FSR, there will be a AM transmission resonance (reflection minimum). If HOMs are also present (created by either partially occluding or misaligning the injection beam), they too will generate transmission resonances, but at a frequency shift proportional to the Gouy phase. For the theoretical 19 deg one-way Gouy phase at the sites, this mode spacing is approximately 300 kHz. If the transmission resonances of two or more modes can be simultaneously measured, their frequency separation will provide a direct measurement of the SRC Gouy phase.
The above figure illustrates this measurement configuration. An attached PDF gives more detail and the expected response based on Finesse modeling of this IFO configuration.
We did a optical measurement of the MCREFL_PD transimpedance using the Jenny Laser set-up. We used 0.56mW @1064nm on the NewFocus 1611 Photodiode as reference and 0.475mW @1064nm on the MCREFL_PD. Transfer function was measured using the AG4395 network analyzer. We also fit the data using the refined LISO model. From the optical measurement, we can see that we do not have a prominent peak at about 300MHz like the one we had from the electrical transimpedence measurement. We also put in the electrical transimpedence measurement as reference. RMS contribution of 300MHz peak to follow.
As per Rana`s advice I have updated the entry with information on the LISO fit quality and parameters used. I have put all the relevant files concerning the above measurement as well as the LISO fit and output files as a zip file "LISO_fit" . I also added a note describing what each file represents. I have also updated the plot with fit parameters and errors as in elog 10406.
I re-installed the MC REFL photodiode. Centered beam on the PD by adjusting steering mirror to maximize the DC signal level (on o'scope) at the DC monitoring port. Curiously, the DC level on the scope (high-Z) was ~2.66V DC, whereas the MEDM screen reports ~twice that value, at ~5.44 "V". We may want to fix this "calibration" (or better yet, use physical units like mW). Noise-eater On/Off comparison of MC error signals to follow.
Till RIN measurement noise eater is off on 2W laser. The inno 1W has no noise eater.
2010 power v current table is below.
Koji and Kevin measured the output power vs injection current for the Innolight 2W laser.
The threshold current is 0.75 A.
The following data was taken with the laser crystal temperature at 25.04ºC (dial setting: 0.12).
MCRefl is absent, it is under investigation. I removed a bunch of hardware and note all spare optics along the edges.
I've removed the MC REFL PD unit from the AS table for investigation. So MC won't lock.
PSL shutter was closed and location of PD was marked with sharpie (placing guides to indicate position wasn't convenient). I also kapton taped the PD to minimize dust settling on the PD while I have it out in the electronics area. Johannes has the camera, and my cellphone image probably isn't really high-res enough for diagnostics but I'm posting it here anyways for what it's worth. More importantly - the board is a D980454 revision B judging by the board, but there is no schematic for this revision on the DCC. The closest I can find is a D980454 Rev D. But I can already see several differences in the component layout (though not all of them may be important). Making a marked up schematic is going to be a pain . I'm also not sure what the specific make of the PD installed is.
The lid of the RF cage wasn't on.
More to follow tomorrow, PD is on the electronics workmench for now...
gautam 28 March 2018: Schematic has been found from secret Dale stash (which exists in addition to the secret Jay stash). It has also been added to the 40m electronics tree.
We made a measurement of the MC_REFL photodiode transfer function using the network analyzer. We did it for two different power input (0dB and -10dB) to the test measurement point of the MC_REFL photodiode. This was important to ensure the measurements of the transfer function of the MC_REFL photodiode was in the linear regime. The measurements are shown in attachment 1. We corrected for phase noise for the length of cable (50cm) used for the measurement. With reference to ELOG 10406, in comparison to the transimpedance measurement performed by Riju and Koji, there is a much stronger peak around 290MHz as observed by our measurement.
We also did a noise measurement for the MC_REFL photodiode. We did it for three scenarios: 1. Without any light falling on the photodiode 2. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was OFF 3. With light falling on the photodiode, the MC misaligned and the NPRO noise eater was ON. We observed that the noise eater does reduce the noise being observed from 80kHz to 20MHz. This is shown in attachment 2.
We did the noise modelling of the MC_REFL photodiode using LISO and tried matching the expected noise from the model with the noise measurements we made earlier. The modeled noise is in good agreement with the measured noise with 10Ohms in the output resistance. The schematic for the MC_REFL photodiode however reveals a 50Ohm resistance being used. The measured noise shows excess noise ~ 290MHz. This is not predicted from the simplied LISO model of the photodiode we took.
Discussion with Koji and Gautam revealed that we do not have the exact circuit diagram for the MC_REFL photodiode. Hence the simplified model that was assumed earlier is not able to predict the excess noise at high frequencies. One thing to note however, is that the excess noise is measured with the same amplitude even with no light falling on the MC_REFL photodiode. This means that there is a positive feedback and oscillation in the op-amp (MAX4107) at high frequencies. One way to refine the LISO model would be to physically examine the photodiode circuit.
We also recorded the POX and POY RF monitor photodiode outputs when the interferometer arms are independently stabilized to the laser. Given the noise outputs from the RF photodiodes were similar, we have only plotted the POY RF monitor output for the sake of clarity and convenience.
While Kevin and Arijit were doing their MC_REFL PD noise measurements (which they will elog about separately shortly), I noticed a feature around 600kHz that reminded me of the NPRO noise eater feature. This is supposed to suppress the relaxation oscillation induced peak in the RIN of the PSL. Surprisingly, the noise eater switch on the NPRO front panel was set to "OFF". Is this the normal operating state? I thought we want the noise eater to be "ON"? Have to measure the RIN on the PSL table itself with one of the many available pick off PDs. In any case, as Attachment #1 showed, turning the noise eater back on did not improve the excess IMC frequency noise.
We setup the channels for PID control of the seismometer can. First, we ssh into c1auxex and went to /cvs/cds/caltech/target/c1auxex2 and found ETMXaux.db. We then added in new soft channels that we named C1:PEM-SEIS_EX_TEMP_SLOWKP, C1:PEM-SEIS_EX_TEMP_SLOWKI, C1:PEM-SEIS_EX_TEMP_SLOWKD that will control the proportional, integral and differential gain respectively. These channels are used in the script FSSSlow.py for PID control. We then had to restart the system, but first we turned off the LSC mode and then shut down the watchdog on the X end. After doing the restart, we disabled the OPLEV as well before restarting the watchdog. Then, we enabled the LSC mode again. This is done to not damage any of the optics during the restart. The restart is done by using sudo systemctl restart modbusIOC.service and restarted with sudo systemctl status modbusIOC.service. Then, we made sure that the channels existed and could be read and writtten to, so we tried z read [channel name] and it read 0.0. We then did z write [channel name] 5, and it wrote 5 to that channel. Now that the channels work, we can implement the PID script and check to make sure that it works as well.
CC1 old MKS cold cathode gauge randomly turns on- off. This makes software interlock close VM1 to protect RGA So the closed off RGA region pressure goes up and the result is distorted RGA scan.
CC1 MKS gauge is disconnected and VM1 opened. This reminds me that we should connect our interlocks to CC1 Hornet Pressure gauge.
Pumpdown 80 at 511 days and pd80b at 218 days
Valve configuration: special vacuum normal, annuloses are not pumped at 3 Torr, IFO pressure 7.4e-6 Torr at vac envelope temp 22 +- 1C degrres